# VGPT-PINN: Viscosity-enhanced Generative Pre-Trained Physics Informed Neural Networks for parameterized nonlinear conservation laws

### v1

@@ Abstract
We propose a Viscosity-enhanced Generative Pre-Trained Physics-Informed Neural Network with a transform layer (VGPT-PINN) for solving parameterized nonlinear conservation laws. The VGPT-PINN extends the traditional physics-informed neural networks and its recently proposed generative pre-trained strategy for linear model reduction to nonlinear model reduction and shock-capturing domains. By utilizing an adaptive meta-network, a simultaneously trained transform layer, viscosity enhancement strategies, implementable shock interaction analysis, and a separable training process, the VGPT-PINN efficiently captures complex parameter-dependent shock formations and interactions. Numerical results of VGPT-PINN applied to the families of inviscid Burgers’ equation and the Euler equations, parameterized by their initial conditions, demonstrate the robustness and accuracy of the proposed technique. It accurately solves for the viscosity solution via very few neurons without leveraging any a priori knowledge of the equations or its initial condition.
@@ 1 Introduction
While being able to accurately capture fully convective phenomena including that of full shocks with varying (linear or simple nonlinear) transport speed, the TGPT-PINN struggles in simulating inviscid fully nonlinear conservation laws. In this paper, we tackle that exact problem by proposing the viscosity-enhanced TGPT-PINN (VGPT-PINN) framework. These problems are challenging due to the emergence of dual challenges, the existence of non-physical solutions and the slow decay of the Kolmogorov $n$-width. In addition to an adaptive meta-network and a simultaneously trained transformation layer that are inherited from the TGPT-PINN, the VGPT-PINN introduces several key innovations to address this dual challenges for machine learning based model order reduction for parametric conservation laws. First, it replaces the original PDE with its characteristic form, simplifying the computation and enabling better capturing of shock waves. To address overfitting near discontinuities, the VGPT-PINN incorporates physics-dependent weights into the loss function, which helps the model focus on regions with shocks and improves solution accuracy. The loss function is further enhanced by including terms for the PDE, initial and boundary conditions, as well as the Rankine-Hugoniot (RH) condition [26, 30], which is essential for accurately modeling physically-relevant shock dynamics. Additionally, an indicator function is used to detect the location of discontinuities, allowing the model to adaptively concentrate on these critical areas and to predict shock intersection time. The VGPT-PINN similarly employs a nonlinear model order reduction approach by introducing a parameter-dependent transform layer, effectively handling parameter-dependent discontinuities and overcoming the limitations of linear model reduction in transport-dominated regimes. Moreover, the method adopts an offline-online structure. During the online stage, only the parameters of the transform layer and output layer are optimized significantly reducing computational effort. In the offline stage, a greedy algorithm is applied, generating a small number of neurons whose activation functions are full PINNs pre-trained at judiciously selected locations in the parameter domain. These inherited features, combined with the above-mentioned innovations, make VGPT-PINN a powerful and efficient tool for solving hyperbolic conservation laws with parameter-dependent discontinuities. To the best of our knowledge, this is the first attempt to accelerate the computing of parametric conservation laws leveraging nonlinear model reduction techniques in the neural network setting.
@@ 2 Background
Here $u(\bm{x},t;\mu)$ represents the solution, $u_{0}(\bm{x},\mu)$ provides the initial value, $\mathcal{F}$ is a differential operator and $\mathcal{G}$ encodes boundary information. Analyzing this system’s behavior and its parameter dependence often requires executing thousands to millions of simulations of the PDE. Simulating it accurately and robustly is typically time-intensive, making the extensive repetition of simulations computationally infeasible with traditional numerical methods. To overcome this challenge, Model Order Reduction (MOR) has been gaining traction in the last two to three decades. The fundamental idea of the projection-based MOR [3] is to construct a low-dimensional approximation of the solution space, and then find as a surrogate solution the Galerkin projection of the high-fidelity solution into this low-dimensional space.
@@
The training process minimizes this loss function through back propagation, incrementally adjusting the network’s weights and biases, toward a better approximation with respect to the underlying physical constraints. PINNs represent a significant advancement in the numerical solution of PDEs thanks to the universal approximation capabilities of DNN [15, 48, 23], existence of software packages [33, 21], and the recent explosion of compute power [1, 45]. However, when it comes to pPDEs, PINNs face the challenge of high computational cost associated with repeated training which is more pronounced than traditional methods due to the more resource-intensive nature of PINNs. To address this, researchers are exploring various strategies, such as transfer learning, where a trained network with one set of parameters is fine-tuned for others [47], and multi-fidelity methods that combine low-fidelity models with high-fidelity PINNs [2, 37]. In [8], the authors introduce a highly reduced neural network named GPT-PINN. This innovative approach reduces the size of PINNs required for unseen parameter values. The GPT-PINN is a network-of-networks, employing pre-trained PINNs as customized activation functions within the neurons of its single hidden layer. Using a mathematically rigorous greedy algorithm with residual-based error indicators, we select parameters $\mu^{1},\mu^{2},\cdots,\mu^{n}$ and obtain $\Psi_{\mathsf{NN}}^{\mu^{1}},\Psi_{\mathsf{NN}}^{\mu^{2}},\cdots,\Psi_{\mathsf{NN}}^{\mu^{n}}$ by training corresponding PINNs. For an untrained parameter $\mu$, we approximate the solution as:
@@ 2.2 Conservative Hyperbolic PDEs
where $\mathbf{U}=(u^{1},u^{2},\cdots,u^{m})^{T}$ denotes the unknown function representing conserved quantities with initial condition $\phi(\bm{x})$ and the specified boundary condition $g(\bm{x},t)$ which can be Dirichlet, Neumann or periodic. Additionally, $\mathbf{F}=(f^{1},f^{2},\cdots,f^{m})^{T}$ is the flux function of the dimensions $m\times d$ of $\mathbf{U}$.
@@
$\mathbf{U}(\bm{x},t)=\mathbf{U}(x-\lambda(\mathbf{U})t,0)=\phi(\bm{x}-\mathbf{F^{\prime}(\bm{U})}t).$ (2.8)
It is easy to see that, if $\mathbf{F^{\prime}(\bm{U})}$ is a constant, (2.7) reduces to the transport equation. The solution manifold, as the constant speed changes, is of rank 1 after an appropriate shifting. This family of pPDEs can then be solved exactly in the model reduction setting by one neuron via a carefully designed algorithm identifying this shift without a priori knowledge of (the constant) $\mathbf{F^{\prime}(\bm{U})}$ [7]. On the other hand, if $\mathbf{F^{\prime}(\bm{U})}=\mathbf{U}$, (2.7) becomes the well-known Burgers’ equation, a fundamental example of nonlinear wave phenomena involving discontinuous solutions. Additionally, the more complex Euler equations express the conservation of mass, momentum and energy, relating the velocity field $u$ and $v$ with the density field $\rho$ and the pressure field $p$[17].
@@
Moreover, by choosing $\mathbf{U}=(\rho,\rho u,\rho v,E)^{T}$ and $\mathbf{F}=\left(\mathbf{F}_{1},\mathbf{F}_{2}\right)=(\rho u,\rho u^{2}+p,\rho uv,u(E+p))^{T},(\rho v,\rho uv,\rho v^{2}+p,v(E+p))^{T})$ in (2.7) with $E=\rho(u^{2}+v^{2})/2+p/(\gamma-1)$, the characteristic form of the 2D Euler equations is:
$\frac{\partial\mathbf{U}}{\partial t}+\mathbf{A}_{1}\frac{\partial\mathbf{U}}{\partial x}+\mathbf{A}_{2}\frac{\partial\mathbf{U}}{\partial y}=0,$
where
$\mathbf{A}_{1}=\mathbf{F}_{1}^{\prime}(\mathbf{U})=\left(\begin{array}[]{rrrr}0&1&0&0\\ (\gamma-1)q-u^{2}&\left(3-\gamma\right)u&(1-\gamma)v&\gamma-1\\ -uv&v&u&0\\ \left((\gamma-1)q-H\right)u&H-(\gamma-1)u^{2}&(1-\gamma)uv&\gamma u\end{array}\right),$
$\mathbf{A}_{2}=\mathbf{F}_{2}^{\prime}(\mathbf{U})=\left(\begin{array}[]{rrrr}0&0&1&0\\ -uv&v&u&0\\ (\gamma-1)q-v^{2}&(1-\gamma)u&\left(3-\gamma\right)v&\gamma-1\\ \left((\gamma-1)q-H\right)v&(1-\gamma)uv&H-(\gamma-1)v^{2}&\gamma v\end{array}\right),$
with $q=(u^{2}+v^{2})/2$ and $H=(E+p)/\rho.$
3 The VGPT-PINN algorithm
In this section, we present the new viscosity-enhanced TGPT-PINN framework, designed to enhance the modeling of parameterized PDEs with complex discontinuities. By reformulating the PDE into its characteristic form and integrating a robust loss function that includes physics-based weights, the Rankine-Hugoniot condition, and an accurate prediction of shock intersection, the full model VGPT-PINN ensures higher fidelity in capturing shock dynamics and discontinuous behavior. A nonlinear model order reduction via a parameter-dependent transform layer addresses parameter-dependent discontinuities effectively, while an offline-online computational structure optimizes efficiency by separating parameterized training and fast inference stages. The following sections detail each of these advancements, demonstrating how they contribute to both the accuracy and computational practicality of VGPT-PINN in solving transport-dominated conservation laws with shocks.
3.1 Physics-informed and viscosity-aware full order model
The solutions of nonlinear conservation laws often develop discontinuities within finite time, even if the initial conditions are smooth [30]. This typically results in a deterioration of solution accuracy near shocks and contact waves [38]. The inherent complexity of handling discontinuities makes these problems particularly challenging to solve by traditional numerical methods [28] warranting special care such as incorporation of artificial viscosity [12, 11, 10] and adoption of slope limiting techniques [13, 50]. There is no exception for network-based approaches.
To make sure that the full PINN achieves reasonable accuracy, we build the loss function based on the weighted characteristic form of the equation. Moreover, we enforce the Rankine–Hugoniot condition to drive the solution to the physically-relevant one during the optimization process, reflected by the $\mathcal{L}_{\rm{RH}}(u)$ term of the loss function:
$\mathcal{L}=\mathcal{L}_{\rm{int}}(u)+\varepsilon_{i}\mathcal{L}_{\rm IC}(u)+\varepsilon_{b}\mathcal{L}_{\rm BC}(u)+\varepsilon_{r}\mathcal{L}_{\rm{RH}}(u).$ (3.1)
@@ 3.1 Physics-informed and viscosity-aware full order model
For the Burgers’ equation, the solution often shows discontinuities with steep gradients or significant differences between neighboring $x$. We focus solely on the $u$ terms, as follows:
@@
Here, $\varepsilon_{1}$ and $\varepsilon_{2}$ are two parameters used to detect jumps in shock waves, and their values can be adjusted depending on the specific problem at hand. Unless otherwise stated, we set $\varepsilon_{1}=\varepsilon_{2}=0.2.$ In the 2D case, the RH relation is coupled with density, velocity and pressure. It can be simplified as [30]:
$\rho_{1}\rho_{2}\left[\left(u_{1}-u_{2}\right)^{2}+\left(v_{1}-v_{2}\right)^{2}\right]=\left(p_{1}-p_{2}\right)\left(\rho_{1}-\rho_{2}\right),$ (3.8)
@@
where subscripts 1 and 2 denote the pre-shock and post-shock states respectively, dependent on $\lambda_{\rm RH}$ and computed by the adjacent points $(\bm{x}\pm\Delta\bm{x},t)$. These two relations also hold in the 1D case by simply taking $v_{1}=v_{2}=0.$ Using the indicator $\lambda_{\rm RH}$ in (3.7), the penalty term for the RH condition in the Euler equations for both 1D and 2D is defined as:
$\mathcal{L}_{\mathrm{RH}}\coloneqq$ $\left\|\lambda_{\rm RH}\left[\rho_{1}\rho_{2}\left(\left(u_{1}-u_{2}\right)^{2}+\left(v_{1}-v_{2}\right)^{2}\right)-\left(p_{1}-p_{2}\right)\left(\rho_{1}-\rho_{2}\right)\right]\right\|^{2}_{2}$ (3.9)
@@ 3.2 Reduced order model with a separable training process
The VGPT-PINN inherits the structure and philosophy of TGPT-PINN while adopting the viscosity-aware loss function (3.1) and a novel separable training. By introducing the transform layer in a GPT-PINN, the VGPT-PINN can capture the parameter-dependent discontinuity locations and achieves good accuracy. The transform layer $\mathcal{T}_{\mu,\eta}$ is designed as a mapping $\mathcal{T}_{\mu,\mu^{i}}(\bm{x},t):\Omega\times[0,T]\longrightarrow\Omega\times[0,T],$ and we employ the linear transformation by Chen et al. [7]
@@
where $\Theta(\mu)\coloneqq\left\{\{W_{\mu,\mu^{i}}\}_{i=1}^{n},\{b_{\mu,\mu^{i}}\}_{i=1}^{n},\{c_{i}(\mu)\}_{i=1}^{n}\right\}$ represents the $n(d^{2}+3d+3)$ parameters to be trained in the VGPT-PINN, and $n$ is the number of snapshots.
@@
The loss function employed in the VGPT-PINN is defined in the same fashion as the full PINN, consisting of the governing equations, the initial value, the boundary condition and the RH condition:
@@
$\mathcal{L}_{\mathrm{int}}^{\mathrm{VGPT}}(\Theta(\mu))=\frac{1}{\left|\mathcal{C}_{o}^{r}\right|}\sum_{(\bm{x},t)\in\mathcal{C}_{o}}\left\|\lambda(\bm{x},t)\,\left(\frac{\partial(\Psi_{\mathrm{\mathsf{NN}}}^{\Theta(\mu)})}{\partial t}+\frac{\partial(\mathbf{F}(\Psi_{\mathrm{\mathsf{NN}}}^{\Theta(\mu)}))}{\partial x}\Psi_{\mathrm{\mathsf{NN}}}^{\Theta(\mu)}\right)\right\|_{2}^{2},$ (3.13a)
@@
$\mathcal{L}_{\mathrm{BC}}^{\mathrm{VGPT}}(\Theta(\mu))=\frac{1}{\mid\mathcal{C}_{\partial}^{r}\mid}\sum_{(\bm{x},t)\in\mathcal{C}_{\partial}}\left\|\mathcal{G}\left(\Psi_{\mathrm{\mathsf{NN}}}^{\Theta(\mu)}\right)(\bm{x},t)\right\|_{2}^{2},$ (3.13b)
@@
$\mathcal{L}_{\mathrm{IC}}^{\mathrm{VGPT}}(\Theta(\mu))=\frac{1}{\left|\mathcal{C}_{i}^{r}\right|}\sum_{\bm{x}\in\mathcal{C}_{i}}\left\|\Psi_{\mathrm{\mathsf{NN}}}^{\Theta(\mu)}(\bm{x},0)-u_{0}(\bm{x})\right\|_{2}^{2},$ (3.13c)
@@
$\mathcal{L}_{\rm{RH}}^{\mathrm{VGPT}}(\Theta(\mu))=\frac{1}{\mid\mathcal{C}_{S}^{r}\mid}\sum_{(\bm{x},t)\in\mathcal{C}_{S}}\left\|\lambda_{{\rm RH},t}\cdot\left(\bm{x}_{{\rm RH},t}+\mathbf{S}_{\rm RH}\cdot{\Delta t}-\bm{x}_{{\rm RH},t+\Delta t}\right)\right\|_{2}^{2}.$ (3.13d)
@@
The total loss function for the VGPT-PINN can be expressed by a combination of all these contributions:
$\mathcal{L}_{\mathrm{PINN}}^{\mathrm{VGPT}}=\mathcal{L}_{\mathrm{int}}^{\mathrm{VGPT}}+\varepsilon_{i}\mathcal{L}_{\mathrm{IC}}^{\mathrm{VGPT}}+\varepsilon_{b}\mathcal{L}_{\mathrm{BC}}^{\mathrm{VGPT}}+\varepsilon_{r}\mathcal{L}_{\rm{RH}}^{\mathrm{VGPT}},$ (3.14)
@@
where $\varepsilon_{i}$, $\varepsilon_{b}$ and $\varepsilon_{r}$ are balancing parameters for the loss terms, aligned with the values in (3.1). By repeatedly updating $\Theta(\mu)$ in the training process, the hyper-reduced network gradually converges towards the target solution. The algorithm details for the VGPT-PINN are presented in Figure 1.
@@
Figure 1: The VGPT-PINN design schematic for Burgers’ (top) and Euler (bottom) equations. For any given parameter value $\mu$, a $\mu$-dependent loss is constructed and the coefficients $\{c_{i}(\mu)\}_{i=1}^{n}$ and the weights and biases in $\{\mathcal{T}_{\mu,\mu^{i}}\}_{i=1}^{n}$ are trained.
@@
The training process focuses on minimizing the loss function $\mathcal{L}_{\mathrm{PINN}}^{\mathrm{VGPT}}$ as defined in (3.13). This is accomplished by utilizing standard techniques such as automatic differentiation and back propagation, employing the same learning rate for all the parameters in $\Theta(\mu)$. By iteratively updating the network’s parameters based on calculated gradients, the VGPT-PINN learns to approximate the desired solution to the equations. However, the existence of discontinuities makes these problems challenging to optimize, especially for the Euler equations.

### v2

# EGPT-PINN: Entropy-enhanced Generative Pre-Trained Physics Informed Neural Networks for parameterized nonlinear conservation laws

@@ Abstract
We propose an entropy-enhanced Generative Pre-Trained Physics-Informed Neural Network with a transform layer (EGPT-PINN) for solving parameterized nonlinear conservation laws. The EGPT-PINN extends the traditional physics-informed neural networks and its recently proposed generative pre-trained strategy for linear model reduction to nonlinear model reduction and shock-capturing domains. By utilizing an adaptive meta-network, a simultaneously trained transform layer, entropy enhancement strategies, implementable shock interaction analysis, and a separable training process, the EGPT-PINN efficiently captures complex parameter-dependent shock formations and interactions. Numerical results of EGPT-PINN applied to the families of inviscid Burgers’ equation and the Euler equations, parameterized by their initial conditions, demonstrate the robustness and accuracy of the proposed technique. It accurately solves the viscosity solution via very few neurons without leveraging any a priori knowledge of the equations or its initial condition. Moreover, via a simple augmentation of the loss function by model-data mismatch, we demonstrate the robustness of EGPT-PINN in solving inverse problems more accurately than the vanilla and entropy-enhanced versions of PINN.
@@ 1 Introduction
While being able to accurately capture fully convective phenomena including that of full shocks with varying (linear or simple nonlinear) transport speed, the TGPT-PINN struggles in simulating inviscid fully nonlinear conservation laws. In this paper, we tackle that exact problem by proposing the Entropy-enhanced TGPT-PINN (EGPT-PINN) framework. These problems are challenging due to the emergence of dual challenges, the existence of non-physical solutions and the slow decay of the Kolmogorov $n$-width. In addition to an adaptive meta-network and a simultaneously trained transformation layer that are inherited from the TGPT-PINN, the EGPT-PINN introduces several key innovations to address this dual challenges for machine learning based model order reduction for parametric conservation laws. First, it replaces the original PDE with its characteristic form, simplifying the computation and enabling better capturing of shock waves. To address overfitting near discontinuities, the EGPT-PINN incorporates physics-dependent weights into the loss function, which helps the model focus on regions with shocks and improves solution accuracy. The loss function is further enhanced by including terms for the PDE, initial and boundary conditions, as well as the Rankine-Hugoniot (RH) condition [26, 30], which is essential for accurately modeling physically-relevant shock dynamics. Additionally, an indicator function is used to detect the location of discontinuities, allowing the model to adaptively concentrate on these critical areas and to predict shock intersection time. The EGPT-PINN similarly employs a nonlinear model order reduction approach by introducing a parameter-dependent transform layer, effectively handling parameter-dependent discontinuities and overcoming the limitations of linear model reduction in transport-dominated regimes. Moreover, the method adopts an offline-online structure. During the online stage, only the parameters of the transform layer and output layer are optimized significantly reducing computational effort. In the offline stage, a greedy algorithm is applied, generating a small number of neurons whose activation functions are full PINNs pre-trained at judiciously selected locations in the parameter domain. These inherited features from TGPT-PINN, combined with the above-mentioned innovations, make EGPT-PINN a powerful and efficient tool for solving hyperbolic conservation laws with parameter-dependent discontinuities. To the best of our knowledge, this is the first attempt to accelerate the computing of parametric conservation laws leveraging nonlinear model reduction techniques in the neural network setting. To further illustrate the robustness of EGPT-PINN, we leverage it for inverse problems. The hallmark feature of PINNs and other neural network-based methods in solving inverse problems is that they do it with a computational cost that is comparable to the forward problem. The main tool is integrating the unknown parameters directly into the training process alongside network parameters. We demonstrate that this advantage translates from the full PINN to EGPT-PINN when solving the challenging boundary inference problem for the Euler equation.
@@ 2 Background
Here $u(\bm{x},t;\mu)$ represents the solution, $u_{0}(\bm{x},\mu)$ provides the initial value, $\mathcal{F}$ is a differential operator and $\mathcal{G}$ encodes boundary information.
Analyzing this system’s behavior and its parameter dependence often requires executing thousands to millions of simulations of the PDE. Simulating it accurately and robustly is typically time-intensive, making the extensive repetition of simulations computationally infeasible with traditional numerical methods. To overcome this challenge, Model Order Reduction (MOR) has been gaining traction in the last few decades. The fundamental idea of the projection-based MOR [3] is to construct a low-dimensional approximation of the solution space, and then find the Galerkin projection of the high-fidelity solution into this low-dimensional space as a surrogate solution.
@@
The training process minimizes this loss function through back propagation, incrementally adjusting the network’s weights and biases, toward a better approximation with respect to the underlying physical constraints. PINNs represent a significant advancement in the numerical solution of PDEs thanks to the universal approximation capabilities of DNN [15, 48, 23], existence of software packages [33, 21], and the recent explosion of compute power [1, 45]. However, when it comes to pPDEs, PINNs face the challenge of high computational cost associated with repeated training which is more pronounced than traditional methods due to the more resource-intensive nature of PINNs.
To address this, researchers are exploring various strategies, such as transfer learning, where a trained network with one set of parameters is fine-tuned for others [47], and multi-fidelity methods that combine low-fidelity models with high-fidelity PINNs [2, 37]. In [8], the authors introduce a highly reduced neural network named GPT-PINN. This innovative approach reduces the size of PINNs required for unseen parameter values. The GPT-PINN is a network-of-networks, employing pre-trained PINNs as customized activation functions within the neurons of its single hidden layer. Using a mathematically rigorous greedy algorithm with residual-based error indicators, we select parameters $\mu^{1},\mu^{2},\cdots,\mu^{n}$ and obtain $\Psi_{\mathsf{NN}}^{\mu^{1}},\Psi_{\mathsf{NN}}^{\mu^{2}},\cdots,\Psi_{\mathsf{NN}}^{\mu^{n}}$ by training corresponding PINNs. For an untrained parameter $\mu$, we approximate the solution as:
@@ 2.2 Conservative Hyperbolic PDEs
where $\mathbf{U}=(u^{1},u^{2},\cdots,u^{m})^{T}$ denotes the unknown function representing conserved quantities with initial condition $\phi(\bm{x})$ and an appropriate boundary condition. Additionally, $\mathbf{F}=(f^{1},f^{2},\cdots,f^{m})^{T}$ is the flux function of the dimensions $m\times d$ of $\mathbf{U}$.
@@
$\mathbf{U}(\bm{x},t)=\mathbf{U}(\bm{x}-\lambda(\mathbf{U})t,0)=\phi(\bm{x}-\mathbf{F^{\prime}(\bm{U})}t).$ (2.8)
It is easy to see that, if $\mathbf{F^{\prime}(\bm{U})}$ is a constant, (2.7) reduces to the transport equation. The solution manifold, as the constant speed changes, is of rank 1 after an appropriate shifting. This family of pPDEs can then be solved exactly in the model reduction setting by one neuron via a carefully designed algorithm identifying this shift without a priori knowledge of (the constant) $\mathbf{F^{\prime}(\bm{U})}$ [7]. On the other hand, if $\mathbf{F^{\prime}(\bm{U})}=\mathbf{U}$, (2.7) becomes the well-known Burgers’ equation, a fundamental example of nonlinear wave phenomena involving discontinuous solutions. Additionally, the more complex Euler equations express the conservation of mass, momentum and energy, relating the velocity field $u$ and $v$ with the density field $\rho$ and the pressure field $p$[17]. Next, we will discuss each of the two equations in detail.
@@
3 The EGPT-PINN algorithm
In this section, we present the new Entropy-enhanced TGPT-PINN framework, designed to enhance the modeling of parameterized PDEs with complex discontinuities. By reformulating the PDE into its characteristic form and integrating a robust loss function that includes physics-based weights, the Rankine-Hugoniot condition, and an accurate prediction of shock intersection, the full model EGPT-PINN ensures higher fidelity in capturing shock dynamics and discontinuous behavior. A nonlinear model order reduction via a parameter-dependent transform layer addresses parameter-dependent discontinuities effectively, while an offline-online computational structure optimizes efficiency by separating parameterized training and fast inference stages. The following sections detail each of these advances and demonstrate how they contribute to both the accuracy and computational practicality of EGPT-PINN in solving transport-dominated conservation laws with shocks.
3.1 Physics-informed and entropy-aware full order model, EPINN
The solutions of nonlinear conservation laws often develop discontinuities within finite time, even if the initial conditions are smooth [30]. This typically results in a deterioration of solution accuracy near shocks and contact waves [38]. The inherent complexity of handling discontinuities makes these problems particularly challenging to solve by traditional numerical methods [28], warranting special care such as incorporation of artificial viscosity [12, 11, 10] and adoption of slope limiting techniques [13, 50]. There is no exception for network-based approaches.
To make sure that the full PINN achieves reasonable accuracy, we build the entropy-aware loss function based on the weighted characteristic form of the equation and enforce the Rankine–Hugoniot condition. To differentiate from the vanilla version, we call the resulting entropy-enhanced PINN EPINN. The loss function of the EPINN drives the solution to the physically-relevant one during the optimization process, reflected by its $\mathcal{L}_{\rm{RH}}(u)$ term:
$\mathcal{L}(u)=\mathcal{L}_{\rm{int}}(u)+\varepsilon_{i}\mathcal{L}_{\rm IC}(u)+\varepsilon_{b}\mathcal{L}_{\rm BC}(u)+\varepsilon_{r}\mathcal{L}_{\rm{RH}}(u).$ (3.1)
@@ 3.1 Physics-informed and entropy-aware full order model, EPINN
For the Burgers’ equation, the solution often shows discontinuities with steep gradients or significant differences between neighboring $\bm{x}$. We focus solely on the $u$ terms, as follows:
@@
Here, $\varepsilon_{1}$ and $\varepsilon_{2}$ are two parameters used to detect jumps in shock waves, and their values can be adjusted depending on the specific problem at hand. Unless otherwise stated, we set $\varepsilon_{1}=\varepsilon_{2}=0.2.$ The RH relation, coupling density velocity and pressure, can be simplified as [30]:
$\rho_{1}\rho_{2}\left(u_{1}-u_{2}\right)^{2}=\left(p_{1}-p_{2}\right)\left(\rho_{1}-\rho_{2}\right),$ (3.8)
@@
where subscripts 1 and 2 denote the pre-shock and post-shock states respectively, dependent on $\lambda_{\rm RH}$ and computed by the adjacent points $(\bm{x}\pm\Delta\bm{x},t)$. Using the indicator $\lambda_{\rm RH}$ in (3.7), the penalty term for the RH condition in the Euler equations is defined as:
$\mathcal{L}_{\mathrm{RH}}\coloneqq$ $\left\|\lambda_{\rm RH}\left[\rho_{1}\rho_{2}\left(u_{1}-u_{2}\right)^{2}-\left(p_{1}-p_{2}\right)\left(\rho_{1}-\rho_{2}\right)\right]\right\|^{2}_{2}$ (3.9)
@@ 3.2 Reduced order model with a separable training process
The EGPT-PINN inherits the structure and philosophy of TGPT-PINN while adopting the entropy-aware loss function (3.1) and a novel separable training. By introducing the transform layer in a GPT-PINN, the EGPT-PINN can capture the parameter-dependent discontinuity locations and achieves good accuracy. The transform layer $\mathcal{T}_{\mu,\eta}$ is designed as a mapping $\mathcal{T}_{\mu,\mu^{i}}(\bm{x},t):\Omega\times[0,T]\longrightarrow\Omega\times[0,T],$ and we employ the linear transformation by Chen et al. [7]
@@
where $\Theta(\mu)\coloneqq\left\{\{W_{\mu,\mu^{i}}\}_{i=1}^{n},\{b_{\mu,\mu^{i}}\}_{i=1}^{n},\{c_{i}(\mu)\}_{i=1}^{n}\right\}$ represents the $n(d^{2}+3d+3)$ parameters to be trained in the EGPT-PINN, and $n$ is the number of snapshots.
@@
The loss function employed in the EGPT-PINN is defined in the same fashion as the full PINN, consisting of the governing equations, the initial value, the boundary condition and the RH condition:
@@
$\mathcal{L}_{\mathrm{int}}^{\mathrm{EGPT}}(\Theta(\mu))=\frac{1}{\left|\mathcal{C}_{o}^{r}\right|}\sum_{(\bm{x},t)\in\mathcal{C}_{o}}\left\|\lambda(\bm{x},t)\,\left(\frac{\partial(\Psi_{\mathrm{\mathsf{NN}}}^{\Theta(\mu)})}{\partial t}+\frac{\partial(\mathbf{F}(\Psi_{\mathrm{\mathsf{NN}}}^{\Theta(\mu)}))}{\partial x}\Psi_{\mathrm{\mathsf{NN}}}^{\Theta(\mu)}\right)\right\|_{2}^{2},$ (3.13a)
@@
$\mathcal{L}_{\mathrm{BC}}^{\mathrm{EGPT}}(\Theta(\mu))=\frac{1}{\mid\mathcal{C}_{\partial}^{r}\mid}\sum_{(\bm{x},t)\in\mathcal{C}_{\partial}}\left\|\mathcal{G}\left(\Psi_{\mathrm{\mathsf{NN}}}^{\Theta(\mu)}\right)(\bm{x},t)\right\|_{2}^{2},$ (3.13b)
@@
$\mathcal{L}_{\mathrm{IC}}^{\mathrm{EGPT}}(\Theta(\mu))=\frac{1}{\left|\mathcal{C}_{i}^{r}\right|}\sum_{\bm{x}\in\mathcal{C}_{i}}\left\|\Psi_{\mathrm{\mathsf{NN}}}^{\Theta(\mu)}(\bm{x},0)-u_{0}(\bm{x})\right\|_{2}^{2},$ (3.13c)
@@
$\mathcal{L}_{\rm{RH}}^{\mathrm{EGPT}}(\Theta(\mu))=\frac{1}{\mid\mathcal{C}_{S}^{r}\mid}\sum_{(\bm{x},t)\in\mathcal{C}_{S}}\left\|\lambda_{{\rm RH},t}\cdot\left(\bm{x}_{{\rm RH},t}+\mathbf{S}_{\rm RH}\cdot{\Delta t}-\bm{x}_{{\rm RH},t+\Delta t}\right)\right\|_{2}^{2}.$ (3.13d)
@@
The total loss function for the EGPT-PINN can be expressed by a combination of all these contributions:
$\mathcal{L}_{\mathrm{PINN}}^{\mathrm{EGPT}}=\mathcal{L}_{\mathrm{int}}^{\mathrm{EGPT}}+\varepsilon_{i}\mathcal{L}_{\mathrm{IC}}^{\mathrm{EGPT}}+\varepsilon_{b}\mathcal{L}_{\mathrm{BC}}^{\mathrm{EGPT}}+\varepsilon_{r}\mathcal{L}_{\rm{RH}}^{\mathrm{EGPT}},$ (3.14)
@@
where $\varepsilon_{i}$, $\varepsilon_{b}$ and $\varepsilon_{r}$ are balancing parameters for the loss terms, aligned with the values in (3.1). By repeatedly updating $\Theta(\mu)$ in the training process, the hyper-reduced network gradually converges towards the target solution. The algorithm details for the EGPT-PINN are presented in Figure 1.
@@
Figure 1: The EGPT-PINN design schematic for Burgers’ (top) and Euler (bottom) equations. For any given parameter value $\mu$, a $\mu$-dependent loss is constructed and the coefficients $\{c_{i}(\mu)\}_{i=1}^{n}$ and the weights and biases in $\{\mathcal{T}_{\mu,\mu^{i}}\}_{i=1}^{n}$ are trained.
@@
The training process focuses on minimizing the loss function $\mathcal{L}_{\mathrm{PINN}}^{\mathrm{EGPT}}$ as defined in (3.13). This is accomplished by utilizing standard techniques such as automatic differentiation and back propagation, employing the same learning rate for all the parameters in $\Theta(\mu)$. By iteratively updating the network’s parameters based on calculated gradients, the EGPT-PINN learns to approximate the desired solution to the equations. However, the existence of discontinuities makes these problems challenging to optimize, especially for the Euler equations.